OpenStudy (anonymous):

88 − 4[−8 + (6 − 10)2]

OpenStudy (anonymous):

152

OpenStudy (anonymous):

z − 31 = 60

OpenStudy (anonymous):

c + 19 = 38

OpenStudy (gw2011):

z-31=60 z=60+31 z=91

OpenStudy (gw2011):

c+19=38 c=38-19 c=19